# Quantum Multiple Eigenvalue Gaussian filtered Search: an efficient and versatile quantum phase estimation method

### 01 Feb 2024 v1

@@
Quantum Multiple Eigenvalue Gaussian filtered Search: an efficient and versatile quantum phase estimation method
Zhiyan Ding Department of Mathematics, University of California, Berkeley Haoya Li Department of Mathematics, Stanford University Lin Lin Department of Mathematics, University of California, Berkeley Applied Mathematics and Computational Research Division, Lawrence Berkeley National Laboratory Challenge Institute for Quantum Computation, University of California, Berkeley HongKang Ni Institute for Computational and Mathematical Engineering, Stanford University Lexing Ying Department of Mathematics, Stanford University Institute for Computational and Mathematical Engineering, Stanford University Ruizhe Zhang Simons Institute for the Theory of Computing, University of California, Berkeley
@@ 1 Introduction
Phase estimation is among the most powerful quantum primitives, offering eigenvalue estimates of a Hamiltonian $H$ when given quantum access to Hamiltonian simulation $\exp(-iHt)$ or block encoding of $H$. While initial algorithms, such as the textbook version of phase estimation [NC00], depend on techniques such as the Quantum Fourier Transform (QFT) and multiple ancilla qubits, recent advances demonstrate that comparable results can be achieved with as few as one ancilla qubit. Despite the much simpler quantum circuit, these “modern” approaches often match or exceed the performance of their predecessors due to enhanced postprocessing techniques. This improvement has been demonstrated in applications such as estimating the ground-state energy, which involves determining the smallest eigenvalue of $H$ and is an important application of phase estimation [AL99, GTC19, LT20, LT22, DLT22, WBC22, ZWJ22, ${}^{+}$23, DL23a, NLY23, ${}^{+}$23]. More recently, this enhancement has also been evident in the broader and more general problem of Multiple Eigenvalue Estimation (MEE) [Som19, SHT22, CG22, ELN22, Som19, DTO22, DL23b, LNY23a, ${}^{+}$23]. A typical example of MEE is estimating the low-lying energies of Hamiltonian $H$, which has many applications, such as determining the electronic and optical properties of materials.
@@
In order to solve MEE for a given quantum Hamiltonian $H\in\mathbb{C}^{M\times M}$, we assume the availability of an initial state |$\ket{\psi}$ that contains several dominant modes. Specifically, let $\left\{(\lambda_{m},\ket{\psi_{m}})\right\}^{M}_{m=1}$ represent pairs of eigenvalues and eigenvectors of $H$. We define $p_{m}=\left|\braket{\psi_{m}}{\psi}\right|^{2}$ as the overlap between the initial state and the $m$-th eigenvector. Our primary assumption in this paper is the Sufficiently Dominant Condition: there exists a set of indices $\mathcal{D}\subset{1,2,\cdots,M}$ such that $p_{\min}=:\min_{i\in\mathcal{D}}p_{i}>p_{\mathrm{tail}}=:\sum_{i\in\mathcal{D}^{c}}p_{i}$, where $\mathcal{D}^{c}=\{1,2,\cdots,M\}\setminus\mathcal{D}$. The eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$ are then called the dominant eigenvalues of $H$ with respect to the initial state |$\ket{\psi}$ (or simply the dominant eigenvalues), and the associated eigenvectors are referred to as the dominant eigenvectors. For simplicity, we assume $\|H\|\leq\pi$, which also implies $\{\lambda_{m}\}_{m\in\mathcal{D}}\subset[-\pi,\pi]$. Our objective is to estimate the dominant eigenvalues $\left\{\lambda_{m}\right\}_{m\in\mathcal{D}}$. From a signal processing perspective, the Sufficiently Dominant Condition allows us to differentiate signal and noise, providing a natural basis for analysis. Different versions of the condition have appeared in previous works [LNY23a, NLY23, DL23b, DL23a, SHT22]. A notable instance is in ground-state energy estimation, where the condition is equivalent to the initial overlap $p_{1}$ between the initial state and the ground state being greater than 0.5. The $\mathcal{O}(1)$ error allowed by this condition is the key to the robustness of many algorithms based on phase estimation.
@@
To estimate the dominant eigenvalues, we assume an oracle access to the Hamiltonian simulation $\exp(-itH)$ for any $t\in\mathbb{R}$. Specifically, given any $t\in\mathbb{R}$, we assume the ability to implement the Hadamard test circuit (see Section 2 for details) to obtain an unbiased estimation to $\braket{\psi}{\exp(-itH)}{\psi}$iiiThroughout this paper, we assume access to $\exp(-itH)$ for any $t\in\mathbb{R}$ for simplicity of presentation. It is straightforward to extend our algorithm to the case with integer powers, where access only to $\exp(-inH)$ for $n\in\mathbb{N}$ is assumed. Refer to Section 5 and Appendix C for details.. Several quantum phase estimation algorithms [LT22, LNY23a, NLY23, DL23b, DL23a, SHT22, ${}^{+}$23] have been developed assuming access to the Hadamard test circuit. In general, these algorithms, including the one proposed in this paper, involve three steps (refer to Fig. 1 for the flowchart): 1. Generate a proper set of $\{t_{n}\}^{N}_{n=1}\subset\mathbb{R}$; 2. Execute the Hadamard test circuits with $t_{n}$ and obtain the dataset $\{(t_{n},Z_{n})\}^{N}_{n=1}$, where $Z_{n}$ is an approximation of $\braket{\psi}{\exp(-it_{n}H)}{\psi}$; 3. Classically post-process $Z_{n}$ to derive the estimation for $\left\{\lambda_{m}\right\}_{m\in\mathcal{D}}$. The efficiency of a quantum phase estimation algorithm is then quantified by two metrics: the maximal runtime denoted by $T_{\max}=\max_{1\leq n\leq N}|t_{n}|$, and the total runtime $T_{\mathrm{total}}=\sum^{N}_{n=1}|t_{n}|$. Here, $T_{\max}$ and $T_{\mathrm{total}}$ approximately measure the depth of the circuit and the total cost of the algorithm, respectively. Although the access to the Hamiltonian simulation $\exp(-itH)$ is assumed to be exact in this work for the sake of simplicity, we anticipate that a certain level of simulation error can be allowed as well by treating the simulation error together with the statistical errors and the effect of non-dominant eigenvectors (see Section 5 for details).
@@
We further define two types of spectral gaps. The first is the spectral gap between the dominant eigenvalues, denoted by $\Delta_{\text{dom}}:=\min_{i,j\in\mathcal{D},j\not=i}\left|\lambda_{i}-\lambda_{j}\right|$. The second is the spectral gap between the dominant eigenvalues and the remaining eigenvalues, denoted by $\Delta:=\min_{i\in\mathcal{D},j\not=i}\left|\lambda_{i}-\lambda_{j}\right|$. In this paper, we use the notations $\mathcal{O}_{[a]}$, $\Omega_{[a]}$, and $\Theta_{[a]}$ to denote that the quantity also polynomially depends on the parameters in $[a]$.
@@
Although previous algorithms may fulfill some of the mentioned properties (see Section 2.1 for a detailed discussion), to our knowledge, our algorithm stands out as the first algorithm that can be proved to simultaneously achieve all four properties, which matches the best available results in the literature. In particular, the “short” circuit depth property is considered important for applications on early fault-tolerant quantum computers [DL23a, KGCJ23].
@@
Property (5) can be satisfied in some of the prior work, such as [${}^{+}$23, DL23a] for ground state energy estimation and [DL23b] for MEE (see the detailed discussion in Section 3). An application of Property (5) is to ensure the algorithm’s efficiency in the presence of global depolarizing noise. In this scenario, it is important to maintain a short depth $T_{\max}=\mathcal{O}(\log(1/\epsilon))$ (setting $\delta=\Theta(\epsilon/\Delta)$) to ensure that $T_{\rm total}$ polynomially depends on $1/\epsilon$ [DDTL23].
@@ 2 Main algorithm and previous work
$G(\theta)=\left|\frac{1}{N}\sum^{N}_{n=1}Z(t_{n})\exp\left(i\theta t_{n}\right)\right|\,.$ (2)
@@
  • To approximate the next dominant eigenvalue, we establish a block interval $\mathcal{I}_{B,1}=[\boldsymbol{\theta}_{1}-d,\boldsymbol{\theta}_{1}+d]$ around $\boldsymbol{\theta}_{1}$ with a suitable $d$. Subsequently, we identify the second maximal point $\boldsymbol{\theta}_{2}$ of $\mathcal{G}(\theta)$ outside the block interval, denoted $\boldsymbol{\theta}_{2}=\operatorname*{arg\,min}_{\theta\in\mathcal{I}^{c}_{B,1}}G(\theta)$.
@@
$a(t)=\left(1-\int^{\sigma T}_{-\sigma T}\frac{1}{\sqrt{2\pi}T}\exp\left(-\frac{s^{2}}{2T^{2}}\right)\textbf{1}_{[-\sigma T,\sigma T]}(s)\mathrm{d}s\right)\delta_{0}(t)+\frac{1}{\sqrt{2\pi}T}\exp\left(-\frac{t^{2}}{2T^{2}}\right)\textbf{1}_{[-\sigma T,\sigma T]}(t)\,.$ (6)
The choice of $a(t)$ is inspired by the fact that the Fourier transform of a Gaussian function remains a Gaussian function and a recent result of spike localization [LNY23b]. The parameter $\sigma$ represents the level of truncation. Specifically, when $\sigma=\infty$, we have $F(x)=\exp\left(-\frac{T^{2}x^{2}}{2}\right)$, reaching its maximum at $x=0$ and exponentially decaying to zero with respect to $T|x|$. Furthermore, the use of the truncated Gaussian ensures that the maximum runtime $T_{\max}=\max_{n}|t_{n}|$ never exceeds $\sigma T$. In the following part of the paper, we may also employ the notation $a_{T}(t)=a(t)$ and $F_{T}(x)=F(x)$ to emphasize the dependence on $T$.
@@
Each evaluation of $X_{n}$ (or $Y_{n}$) only requires running the Hadamard test circuit with $W=I$ (or $W=S^{\dagger}$) once at $t=t_{n}$. Referring to (7), we obtain:
$\mathbb{E}(Z_{n})=\left\langle\psi\right|\exp(-it_{n}H)\ket{\psi},\quad|Z_{n}|\leq\sqrt{2}.$ (9)
@@
    which poses an undesirable requirement in the context of quantum phase estimation. Here, the last equality is based on the assumption that all the eigenvalues of $H$ belong to the interval $[-\pi,\pi]$.
@@
In Section 3, we present several results that demonstrate that our algorithm can autonomously address the two concerns mentioned above. Specifically, we will show that under condition $p_{\min}>p_{\mathrm{tail}}$, even if $T$ does not satisfy (11), and (5) does not hold, the set $\{\boldsymbol{\theta}_{k}\}^{|\mathcal{D}|}_{k=1}$ still serves as an approximation to the set of dominant eigenvalues in a meaningful sense. Moreover, even when $G(\theta)$ is not $\mathcal{O}(\epsilon)$ close to $\mathcal{G}(\theta)$, the search process remains stable enough to ensure the accuracy of the approximation.
@@ 2.1 Comparison with previous work
The first work using the Hadamard test circuit to achieve Heisenberg-limited scaling is [LT22]. Several works have been developed to achieve Heisenberg-limited scaling for multiple eigenvalue estimation and without relying on any gap assumptions (Properties (2) and (3)iiiiiiRigorously speaking, Heisenberg-limited scaling without gap requirement surpasses the strength of Properties (2) and (3) stated in Section 1. It implies that the algorithm can achieve Heisenberg-limited scaling without any gap assumption. Specifically, for any $\epsilon>0$, the algorithm achieves $\epsilon$-accuracy with $T_{\mathrm{total}}=\widetilde{\mathcal{O}}_{p_{\mathrm{tail}},p_{\min}}(1/\epsilon)$, where the constant is independent of $\Delta,\Delta_{\text{dom}}$. As demonstrated in Theorem 3.1, our algorithm achieves Heisenberg-limited scaling without relying on any gap assumptions.). For example, [DTO22] introduced a method for estimating multiple eigenvalue phases that extends the idea of robust phase estimation (RPE) [${}^{+}$09, KLY15, NLY23] to multiple eigenvalues and achieves Heisenberg-limited scaling without relying on any assumptions about the spectral gap. However, the theoretical analysis in [DTO22] is based on the assumption that all non-dominant modes vanish, as defined in [DTO22, Definition 3.1 and Theorem 4.5], specifically $p_{\mathrm{tail}}=0$. This drawback has been resolved by a recent work [LNY23a]. The robust multiple phase estimation (RMPE) method in [LNY23a] extends the Kitaev-type RPE method [KLY15, NLY23] to the multiple eigenvalue estimation problem by using adaptive simulation time amplifying factors and suitable signal processing algorithms. In contrast to [DTO22], the algorithm proposed in [LNY23a, Section III] can achieve Heisenberg-limited scaling without the requirement of a gap between dominant eigenvalues, even when dealing with an imperfect initial state ($p_{\mathrm{tail}}>0$) (Properties (1)-(3)). The design of the algorithm is based on a dedicated line spectrum estimation algorithm [LNY23b], which employs the same Gaussian filtering function as employed in this work. Moreover, in scenarios where a gap exists between dominant eigenvalues, the algorithm presented in [LNY23a, Section V] combines the line spectrum estimation algorithm with ESPRIT to estimate the dominant eigenvalues with a short circuit depth (Property (4)). Specifically, to achieve $\epsilon$-accuracy, the algorithm requires $T_{\max}=\mathcal{O}_{p_{\min},|\mathcal{D}|}(\max\{p_{\mathrm{tail}}^{-1}\Delta_{\text{dom}}^{-1},p_{\mathrm{tail}}\epsilon^{-1}\})$ and $T_{\mathrm{total}}=\mathcal{O}_{p_{\min},|\mathcal{D}|}(p_{\mathrm{tail}}^{-2}\Delta_{\text{dom}}^{-1}+p_{\mathrm{tail}}^{-1}\epsilon^{-1})$, resulting in a short circuit depth as long as $\epsilon=\mathcal{O}(p_{\mathrm{tail}}^{2}\Delta_{\text{dom}})$. Compared to the algorithms outlined in [LNY23a, Sections III, V], the quantum cost of QMEGS exhibits a logarithmic dependence on $|\mathcal{D}|$, and QMEGS can further achieve the “constant” depth property in the presence of a spectral gap between dominant and other eigenvalues. In addition, QMEGS does not require meticulous parameter adjustment (e.g., $\sigma,T,N,\alpha,q$) based on prior knowledge of $\{(p_{m},\lambda_{m})\}$, and can be more flexible than the algorithms proposed in [LNY23a].
@@
An optimization-based signal processing method called the multi-modal, multi-level quantum complex exponential least squares (MM-QCELS) method has recently been proposed  [DL23b], which generalizes the quantum complex exponential least squares (QCELS) method [DL23a] to the setting of multiple eigenvalues. When $\Delta_{\text{dom}}>0$, MM-QCELS can approximate these dominant eigenvalues with Heisenberg limited scaling, “short”, or “constant” circuit depth (Properties (1), (2), (4), (5)). Furthermore, the quantum cost of MM-QCELS only depends logarithmically on $|\mathcal{D}|$ (Property (6)). To achieve Heisenberg-limited scaling, MM-QCELS needs to generate a sequence of datasets on quantum computers. The classical optimization procedure of MM-QCELS to find $K$ dominant eigenvalues solves an optimization problem in a $K$-dimensional space (see Appendix B for a brief overview). In the worst-case scenario, the classical post-processing cost can grow exponentially in $\epsilon^{-K}$. Compared to MM-QCELS, QMEGS has a much simpler data generation and searching process. The algorithm only requires a single dataset generated by a single $T$. Additionally, leveraging block intervals in the searching process, QMEGS can achieve $\epsilon$-accuracy with Heisenberg-limited scaling even when $\epsilon=\Omega(\Delta_{\text{dom}})$. Regarding the classical processing cost, QMEGS only requires evaluating the filter function at a finite number of discrete points in $[-\pi,\pi]$ and the number of evaluations is $\mathcal{O}(1/\epsilon)$ and is independent of $K$.

### 26 Sep 2024 v2

@@
Abstract
@@ 1 Introduction
Phase estimation is among the most powerful quantum primitives, offering eigenvalue estimates of a Hamiltonian $H$ when given quantum access to Hamiltonian simulation $\exp(-iHt)$ or block encoding of $H$. While initial algorithms, such as the textbook version of phase estimation [1], depend on techniques such as the Quantum Fourier Transform (QFT) and multiple ancilla qubits, recent advances demonstrate that comparable results can be achieved with as few as one ancilla qubit. Despite the much simpler quantum circuit, these “modern” approaches often match or exceed the performance of their predecessors due to enhanced postprocessing techniques. This improvement has been demonstrated in applications such as estimating the ground-state energy, which involves determining the smallest eigenvalue of $H$ and is an important application of phase estimation [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. More recently, this enhancement has also been evident in the broader and more general problem of Multiple Eigenvalue Estimation (MEE) [13, 14, 15, 16, 17, 18, 19, 20]. A typical example of MEE is estimating the low-lying energies of Hamiltonian $H$, which has many applications, such as determining the electronic and optical properties of materials.
@@
In order to solve MEE for a given quantum Hamiltonian $H\in\mathbb{C}^{M\times M}$, we assume the availability of an initial state |$\ket{\psi}$ that contains several dominant modes. Specifically, let $\left\{(\lambda_{m},\ket{\psi_{m}})\right\}^{M}_{m=1}$ represent pairs of eigenvalues and eigenvectors of $H$. We define $p_{m}=\left|\braket{\psi_{m}}{\psi}\right|^{2}$ as the overlap between the initial state and the $m$-th eigenvector. Our primary assumption in this paper is the Sufficiently Dominant Condition: there exists a set of indices $\mathcal{D}\subset{1,2,\cdots,M}$ such that $p_{\min}=:\min_{i\in\mathcal{D}}p_{i}>p_{\mathrm{tail}}=:\sum_{i\in\mathcal{D}^{c}}p_{i}$, where $\mathcal{D}^{c}=\{1,2,\cdots,M\}\setminus\mathcal{D}$. The eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$ are then called the dominant eigenvalues of $H$ with respect to the initial state |$\ket{\psi}$ (or simply the dominant eigenvalues), and the associated eigenvectors are referred to as the dominant eigenvectors. For simplicity, we assume $\|H\|\leq\pi$, which also implies $\{\lambda_{m}\}_{m\in\mathcal{D}}\subset[-\pi,\pi]$. Our objective is to estimate the dominant eigenvalues $\left\{\lambda_{m}\right\}_{m\in\mathcal{D}}$. From a signal processing perspective, the Sufficiently Dominant Condition allows us to differentiate signal and noise, providing a natural basis for analysis. Different versions of the condition have appeared in previous works [19, 11, 18, 10, 14]. A notable instance is in ground-state energy estimation, where the condition is equivalent to the initial overlap $p_{1}$ between the initial state and the ground state being greater than 0.5. The $\mathcal{O}(1)$ error allowed by this condition is the key to the robustness of many algorithms based on phase estimation.
@@
To estimate the dominant eigenvalues, we assume an oracle access to the Hamiltonian simulation $\exp(-itH)$ for any $t\in\mathbb{R}$. Specifically, given any $t\in\mathbb{R}$, we assume the ability to implement the Hadamard test circuit (see Section 2 for details) to obtain an unbiased estimation to $\braket{\psi}{\exp(-itH)}{\psi}$iiiThroughout this paper, we assume access to $\exp(-itH)$ for any $t\in\mathbb{R}$ for simplicity of presentation. It is straightforward to extend our algorithm to the case with integer powers, where access only to $\exp(-inH)$ for $n\in\mathbb{N}$ is assumed. Refer to Section 5 and Appendix C for details.. Several quantum phase estimation algorithms [5, 19, 11, 18, 10, 14, 9] have been developed assuming access to the Hadamard test circuit. In general, these algorithms, including the one proposed in this paper, involve three steps (refer to Fig. 1 for the flowchart): 1. Generate a proper set of $\{t_{n}\}^{N}_{n=1}\subset\mathbb{R}$; 2. Execute the Hadamard test circuits with $t_{n}$ and obtain the dataset $\{(t_{n},Z_{n})\}^{N}_{n=1}$, where $Z_{n}$ is an approximation of $\braket{\psi}{\exp(-it_{n}H)}{\psi}$; 3. Classically post-process $Z_{n}$ to derive the estimation for $\left\{\lambda_{m}\right\}_{m\in\mathcal{D}}$. The efficiency of a quantum phase estimation algorithm is then quantified by two metrics: the maximal runtime denoted by $T_{\max}=\max_{1\leq n\leq N}|t_{n}|$, and the total runtime $T_{\mathrm{total}}=\sum^{N}_{n=1}|t_{n}|$. Here, $T_{\max}$ and $T_{\mathrm{total}}$ approximately measure the depth of the circuit and the total cost of the algorithm, respectively. Although the access to the Hamiltonian simulation $\exp(-itH)$ is assumed to be exact in this work for the sake of simplicity, we anticipate that a certain level of simulation error can be allowed as well by treating the simulation error together with the statistical errors and the effect of non-dominant eigenvectors (see Section 5 for details).
@@
We further define two types of spectral gaps. The first is the spectral gap between the dominant eigenvalues, denoted by $\Delta_{\text{dom}}:=\min_{i,j\in\mathcal{D},j\not=i}\left|\lambda_{i}-\lambda_{j}\right|$. The second is the spectral gap between the dominant eigenvalues and the remaining eigenvalues, denoted by $\Delta:=\min_{i\in\mathcal{D},j\not=i}\left|\lambda_{i}-\lambda_{j}\right|$. In this paper, we use the notations $\mathcal{O}_{[a]}$, $\Omega_{[a]}$, and $\Theta_{[a]}$ to indicate that the quantity polynomially depends on the parameters in $[a]$ if $a$ is a number greater than one, or inversely proportional to $[a]$ if $a$ is a number smaller than one.
@@
Although previous algorithms may fulfill some of the mentioned properties (see Section 2.1 for a detailed discussion), to our knowledge, our algorithm stands out as the first algorithm that can be rigorously proven to simultaneously achieve all four properties, which matches the best available results in the literature. In particular, the “short” circuit depth property is considered important for applications on early fault-tolerant quantum computers [10, 21].
@@
Property (5) can be satisfied in some of the prior work, such as [9, 10] for ground state energy estimation and [18] for MEE (see the detailed discussion in Section 3). An application of Property (5) is to ensure the algorithm’s efficiency in the presence of global depolarizing noise. In this scenario, it is important to maintain a short depth $T_{\max}=\mathcal{O}(\log(1/\epsilon))$ (setting $\delta=\Theta(\epsilon/\Delta)$) to ensure that $T_{\rm total}$ polynomially depends on $1/\epsilon$ [22].
@@ 2 Main algorithm and previous work
$G(\theta)=\left|\frac{1}{N}\sum^{N}_{n=1}Z(t_{n})\exp\left(i\theta t_{n}\right)\right|\,,$ (2)
where $Z(t_{n})$ is a random variable that can take values in {±$\{\pm 1\pm i\}$; a detailed discussion can be found later.
@@
  • To approximate the next dominant eigenvalue, we establish a block interval $\mathcal{I}_{B,1}=[\boldsymbol{\theta}_{1}-d,\boldsymbol{\theta}_{1}+d]$ around $\boldsymbol{\theta}_{1}$ with a suitable $d$. Subsequently, we identify the second maximal point $\boldsymbol{\theta}_{2}$ of $\mathcal{G}(\theta)$ outside the block interval, denoted $\boldsymbol{\theta}_{2}=\operatorname*{arg\,max}_{\theta\in\mathcal{I}^{c}_{B,1}}G(\theta)$.
@@
$a(t)=$ $\left(1-\int^{\sigma T}_{-\sigma T}\frac{1}{\sqrt{2\pi}T}\exp\left(-\frac{s^{2}}{2T^{2}}\right)\textbf{1}_{[-\sigma T,\sigma T]}(s)\mathrm{d}s\right)\delta_{0}(t)$ (6)
$+\frac{1}{\sqrt{2\pi}T}\exp\left(-\frac{t^{2}}{2T^{2}}\right)\textbf{1}_{[-\sigma T,\sigma T]}(t)\,.$
Here, $\delta_{0}(t)$ is Dirac delta function at point 00. The choice of $a(t)$ is inspired by the fact that the Fourier transform of a Gaussian function remains a Gaussian function and a recent result of spike localization [23]. The parameter $\sigma$ represents the level of truncation. Specifically, when $\sigma=\infty$, we have $F(x)=\exp\left(-\frac{T^{2}x^{2}}{2}\right)$, reaching its maximum at $x=0$ and exponentially decaying to zero with respect to $T|x|$. Furthermore, the use of the truncated Gaussian ensures that the maximum runtime $T_{\max}=\max_{n}|t_{n}|$ never exceeds $\sigma T$. In the following part of the paper, we may also employ the notation $a_{T}(t)=a(t)$ and $F_{T}(x)=F(x)$ to emphasize the dependence on $T$.
@@
Here, $X_{n}$ and $Y_{n}$ are random variables that each take values of $-1$ or $1$. Each evaluation of $X_{n}$ (or $Y_{n}$) only requires running the Hadamard test circuit with $W=I$ (or $W=S^{\dagger}$) once at $t=t_{n}$. Referring to (7), we obtain:
$\mathbb{E}(Z_{n})=\left\langle\psi\right|\exp(-it_{n}H)\ket{\psi},\quad Z_{n}\in\left\{\pm 1\pm i\right\},\quad|Z_{n}|=\sqrt{2}.$ (9)
@@
    which poses an undesirable requirement in the context of quantum phase estimationiiiiiiFor example, in ground state energy estimation, it is well known that $T_{\max}$ can be independent of the spectral gap $\lambda_{2}-\lambda_{1}$.. Here, the last equality is based on the assumption that all the eigenvalues of $H$ belong to the interval $[-\pi,\pi]$.
@@
In Section 3, we present several results that demonstrate that our algorithm can autonomously address the two concerns mentioned above. We will show that under condition $p_{\min}>p_{\mathrm{tail}}$, (5) and (11) are not necessary. Specifically, even if $T$ does not satisfy (11), and (5) does not hold, the set $\{\boldsymbol{\theta}_{k}\}^{|\mathcal{D}|}_{k=1}$ still serves as an approximation to the set of dominant eigenvalues in a meaningful sense. Moreover, even when $G(\theta)$ is not $\mathcal{O}(\epsilon)$ close to $\mathcal{G}(\theta)$, the search process remains stable enough to ensure the accuracy of the approximation.
@@ 2.1 Comparison with previous work
The first work using the Hadamard test circuit to achieve Heisenberg-limited scaling is [5]. Several works have been developed to achieve Heisenberg-limited scaling for multiple eigenvalue estimation and without relying on any gap assumptions (Properties (2) and (3)iiiiiiiiiRigorously speaking, Heisenberg-limited scaling without gap requirement surpasses the strength of Properties (2) and (3) stated in Section 1. It implies that the algorithm can achieve Heisenberg-limited scaling without any gap assumption. Specifically, for any $\epsilon>0$, the algorithm achieves $\epsilon$-accuracy with $T_{\mathrm{total}}=\widetilde{\mathcal{O}}_{p_{\mathrm{tail}},p_{\min}}(1/\epsilon)$, where the constant is independent of $\Delta,\Delta_{\text{dom}}$. As demonstrated in Theorem 3.1, our algorithm achieves Heisenberg-limited scaling without relying on any gap assumptions.). For example, [17] introduced a method for estimating multiple eigenvalue phases that extends the idea of robust phase estimation (RPE) [24, 25, 11] to multiple eigenvalues and achieves Heisenberg-limited scaling without relying on any assumptions about the spectral gap. However, the theoretical analysis in [17] is based on the assumption that all non-dominant modes vanish, as defined in [17, Definition 3.1 and Theorem 4.5], specifically $p_{\mathrm{tail}}=0$. This drawback has been resolved by a recent work [19]. The robust multiple phase estimation (RMPE) method in [19] extends the Kitaev-type RPE method [25, 11] to the multiple eigenvalue estimation problem by using adaptive simulation time amplifying factors and suitable signal processing algorithms. In contrast to [17], the algorithm proposed in [19, Section III] can achieve Heisenberg-limited scaling without the requirement of a gap between dominant eigenvalues, even when dealing with an imperfect initial state ($p_{\mathrm{tail}}>0$) (Properties (1)-(3)). The design of the algorithm is based on a dedicated line spectrum estimation algorithm [23], which employs the same Gaussian filtering function as employed in this work. Moreover, in scenarios where a gap exists between dominant eigenvalues, the algorithm presented in [19, Section V] combines the line spectrum estimation algorithm with ESPRIT to estimate the dominant eigenvalues with a short circuit depth (Property (4)). Specifically, to achieve $\epsilon$-accuracy, the algorithm requires $T_{\max}=\mathcal{O}_{p_{\min},|\mathcal{D}|}(\max\{p_{\mathrm{tail}}^{-1}\Delta_{\text{dom}}^{-1},p_{\mathrm{tail}}\epsilon^{-1}\})$ and $T_{\mathrm{total}}=\mathcal{O}_{p_{\min},|\mathcal{D}|}(p_{\mathrm{tail}}^{-2}\Delta_{\text{dom}}^{-1}+p_{\mathrm{tail}}^{-1}\epsilon^{-1})$, resulting in a short circuit depth as long as $\epsilon=\mathcal{O}(p_{\mathrm{tail}}^{2}\Delta_{\text{dom}})$. Compared to the algorithms outlined in [19, Sections III, V], the quantum cost of QMEGS exhibits a logarithmic dependence on $|\mathcal{D}|$, and QMEGS can further achieve the “constant” depth property in the presence of a spectral gap between dominant and other eigenvalues. In addition, QMEGS does not require meticulous parameter adjustment (e.g., $\sigma,T,N,\alpha,q$) based on prior knowledge of $\{(p_{m},\lambda_{m})\}$, and can be more flexible than the algorithms proposed in [19].
@@
An optimization-based signal processing method called the multi-modal, multi-level quantum complex exponential least squares (MM-QCELS) method has recently been proposed  [18], which generalizes the quantum complex exponential least squares (QCELS) method [10] to the setting of multiple eigenvalues. When $\Delta_{\text{dom}}>0$, MM-QCELS can approximate these dominant eigenvalues with Heisenberg limited scaling, “short”, or “constant” circuit depth (Properties (1), (2), (4), (5)). Furthermore, the quantum cost of MM-QCELS only depends logarithmically on $|\mathcal{D}|$ (Property (6)). To achieve Heisenberg-limited scaling, MM-QCELS needs to generate a sequence of datasets on quantum computers. The classical optimization procedure of MM-QCELS to find $K$ dominant eigenvalues solves an optimization problem in a $K$-dimensional space (see Appendix B for a brief overview). In the worst-case scenario, the classical post-processing cost can grow exponentially in $\epsilon^{-K}$. Compared to MM-QCELS, QMEGS has a much simpler data generation and searching process. The algorithm only requires a single dataset generated by a single $T$. Additionally, leveraging block intervals in the searching process, QMEGS can achieve $\epsilon$-accuracy with Heisenberg-limited scaling even when $\epsilon=\Omega(\Delta_{\text{dom}})$. Regarding the classical processing cost, QMEGS only requires evaluating the filter function at a finite number of discrete points in $[-\pi,\pi]$ and the number of evaluations is $\mathcal{O}(1/\epsilon)$ and is independent of $K$.